# GT-Loc: Unifying When and Where in Images Through a Joint Embedding Space

### v1

@@ 2 Related Work
Global geo-localization: Geo-localization, the task of estimating the geographic coordinates of an image, has gained substantial popularity in recent years. Traditionally, geo-localization methods have adopted either a classification approach [40, 27, 37, 19, 22] or an image retrieval approach [24, 28, 29, 34, 48, 49]. The classification approach divides the Earth into a fixed number of geo-cells, assigning the center coordinate of the selected class as the GPS prediction. However, this can result in significant errors depending on the size of the geo-cells, even when the correct class is selected. In contrast, the image retrieval approach compares a query image to a gallery and retrieves the image with the highest similarity. GeoCLIP [36] addresses the limitations of traditional approaches by framing global geo-localization as a GPS retrieval problem. It leverages the pretrained CLIP [23] ViT and employs contrastive learning to align image and location embeddings in a shared feature space. Other methods, such as PIGEON [5], use a hybrid strategy: first using image classification to identify the top-$k$ geo-cells with the highest probability, followed by a secondary retrieval stage for refinement within and across geo-cells. However, PIGEON’s dependence on additional metadata for training—such as administrative boundaries, climate, and traffic—poses a significant limitation. Finally, recent methods such as Img2Loc [46] have begun leveraging multimodal large language models (MLLMs) and retrieval-augmented generation (RAG) to achieve competitive geo-localization performance without the need for dedicated training. However, the effectiveness of these methods is highly dependent on the underlying MLLM and results in substantial inference overhead.
@@ 3 Method
Given a training dataset $S_{train}=\{(I_{i},G_{i},D_{i})\}_{i=1}^{N}$ consisting of image $I_{i}$, GPS coordinates $G_{i}$ and date-time $D_{i}$ triplets, our objective is to train a model that can simultaneously predict the location, time-of-day (ToD) and time-of-year (ToY) from unseen images. Our GT-Loc method consists of three encoders: Image Encoder ($\mathcal{V}$), Location Encoder ($\mathscr{L}$), and Time Encoder ($\mathcal{T}$) as shown in Figure 2. Both geo-localization and time prediction are framed as a retrieval problem. Given a query image $I^{Q}\in S_{eval}$, we compute an image embedding, $V^{Q}=\mathcal{V}(I^{Q})$, using a pre-trained Vision Transformer. Similarly, given a gallery of latitude-longitude pairs, and a gallery of timestamps, we respectively compute galleries of location embeddings $L^{Q}_{k}=\mathcal{L}(G^{Q}_{k})$ and time embeddings $T^{Q}_{k}=\mathcal{T}(D^{Q}_{k})$. In order to predict the location and time, the image embedding is compared against both galleries. The GPS and timestamp with the highest cosine similarity to the image are selected as predictions.
@@ 3.2 Time Representation
The capture time of an image is usually a Unix timestamp, an integer tracking the seconds (or milliseconds) elapsed since January 1, 1970. We discard the year information from the timestamp, as predicting the year is beyond the scope of this work, and instead focus on Time-of-Year (ToY; i.e., month) and Time-of-Day prediction (ToD; i.e., hour). Both ToY and ToD are cyclical, with periods of 12 months and 24 hours, respectively. To convert the Unix timestamp $U_{i}$ into a time representation that focuses on the months and hours, we transform it into a date tuple $D_{i}=\mathrm{\texttt{unix2tuple}}(U_{i})=(m_{i},d_{i},H_{i},M_{i},S_{i})$ with the month, day, hour, minute, and second. From this tuple, we then compute a new time representation composed of the normalized cyclic month-hour pair $D_{i}^{\prime}=(\theta_{i},\phi_{i})$ using
@@ 3.3 Temporal Metric Learning
Since $\mathbf{p}_{i}$ and $\mathbf{q}_{i}$ are normalized, they essentially represent probability mass functions. Thus, we define the image-time contrastive loss by computing the cross-entropy ($\mathrm{CE}$) between $\mathbf{p}_{i}$ and $\mathbf{q}_{i}$:
@@ 4 Experiments
Table 2: Geo-localization results on Im2GPS3k & GWS15k datasets, reported on the ratio of samples that are correctly predicted under distance thresholds of 1 km radius.
@@ 4.2 Ablation Studies
In this section, we present additional experiments to evaluate the effectiveness of our proposed Temporal Metric Loss. Furthermore, in the supplementary material (Section 13), we explore alternative image backbones (i.e., DINOv2 [20], OpenCLIP Ilharco et al. [7]), different time encoders (i.e., Time2Vec Kazemi et al. [10], Circular Decomposition Mac Aodha et al. [16]), and various temporal resolutions to provide additional insights and justification for our design choices.
@@ 4.4 Compositional Image Retrieval
In this section, we evaluate GT-Loc’s capability for compositional image retrieval tasks, leveraging its unified multimodal embedding space. Specifically, we explore retrieving images based on joint queries consisting of both location and time information. Given query location $L^{Q}$ and time $T^{Q}$, we generate a multimodal embedding by averaging their embeddings, inspired by multimodal retrieval methods proposed in prior works [30, 31].
@@ 13.2 Time representation
Motivated by the cyclical nature of time, Mac Aodha et al. 16 used circular decomposition to wrap the temporal input to their geographical prior encoder, resulting in similar embeddings for dates that are close to the start and end of the year, such as December $31^{st}$ and January $1^{st}$. To achieve this, for each dimension $l$ of the temporal input $\vec{x}$, they perform the mapping $[\sin(\pi x^{l}),\cos(\pi x^{l})]$, resulting in two numbers for each dimension.
@@ 14 Compositional image retrieval details
To adapt it for compositional image retrieval, we take a query time $T^{Q}$ and location $L^{Q}$ along with an image gallery ${I^{G}}$. First, we concatenate each image embedding $C_{I}(I^{G})$ with $C_{T}(T^{Q})$ or $C_{L}(L^{Q})$ and feed these pairs to the respective classification heads. This yields probability distributions $P_{i}(l\mid C_{I},C_{T})$ and $P_{i}(t\mid C_{I},C_{L})$ for each image $i$ in the gallery, indicating how well that image matches the queried location and time. Since we already know the desired location and time, we extract the corresponding probabilities from the distributions and average them to produce a final similarity score. Ranking by this score allows us to retrieve the top-$k$ gallery images most likely to match $(T^{Q},L^{Q})$. This adaptation of the Zhai et al. [44] model provides a direct, fair comparison to GT-Loc’s performance on compositional retrieval tasks.
@@ 21 Limitations and reproducibility challenges of existing time prediction methods
Most previous time prediction methods suffer from a lack of standardization in their training and evaluation protocols. For instance, Zhai et al. [44] use subsets of the AMOS [8, 9] and YFCC100M [33] datasets, without providing the source code or exact dataset splits necessary to replicate their experiments. Additionally, their time prediction evaluation relies on cumulative error plots, but they do not provide a single numerical value summarizing the performance of their model. Similarly, while Salem et al. [25] and Padilha et al. [21] offer datasets and code, they do not include the cross-camera split for zero-shot time prediction—a more challenging and informative evaluation protocol that we adopt. Moreover, their results are presented only qualitatively, though they can be adapted for obtaining quantitative results. Salem et al. [26] also omit critical details such as dataset splits for the SkyFinder dataset, do not clarify whether their results corresponds to same- or cross-camera evaluation, and fail to provide the source code for replication. Their use of top-$k$ accuracy as an evaluation metric further complicates direct comparisons. Notably, none of these methods, with the exception of Salem et al. [25] and Padilha et al. [21], compare their time prediction performances against each other, and even these comparisons are only qualitative. Other time prediction approaches, including those by Tsai et al. [35], Li et al. [15], Lalonde et al. [13], and Hold-Geoffroy et al. [6], also face similar challenges, such as a lack of available source code, missing datasets, or datasets that are no longer hosted online. This lack of standardization prevents consistent benchmarking across different methods.

### v2

@@ 2 Related Work
[…]
@@ 3 Method
Given a training dataset $S_{train}=\{(I_{i},G_{i},D_{i})\}_{i=1}^{N}$ consisting of image $I_{i}$, GPS coordinates $G_{i}$ and date-time $D_{i}$ triplets, our objective is to train a model that can simultaneously predict the location, time-of-day (ToD) and time-of-year (ToY) from unseen images. Our GT-Loc method consists of three encoders: Image Encoder ($\mathcal{V}$), Location Encoder ($\mathscr{L}$), and Time Encoder ($\mathcal{T}$) as shown in Figure 2. Both geo-localization and time prediction are framed as a retrieval problem. Given a query image $I^{Q}\in S_{eval}$, we compute an image embedding, $V^{Q}=\mathcal{V}(I^{Q})$, using a pre-trained Vision Transformer. Similarly, given a gallery of latitude-longitude pairs, and a gallery of timestamps, we respectively compute galleries of location embeddings $L^{Q}_{k}=\mathscr{L}(G^{Q}_{k})$ and time embeddings $T^{Q}_{k}=\mathcal{T}(D^{Q}_{k})$. In order to predict the location and time, the image embedding is compared against both galleries. The GPS and timestamp with the highest cosine similarity to the image are selected as predictions.
@@ 3.2 Time Representation
The capture time of an image is usually a Unix timestamp, an integer tracking the seconds (or milliseconds) elapsed since January 1, 1970. We discard the year information from the timestamp, as predicting the year is beyond the scope of this work, and instead focus on Time-of-Year (ToY; i.e., month) and Time-of-Day prediction (ToD; i.e., hour). Both ToY and ToD are cyclical, with periods of 12 months and 24 hours, respectively. To convert the Unix timestamp $U_{i}$ into a time representation that focuses on the months and hours, we transform it into a date-time tuple $D_{i}=\mathrm{\texttt{unix2tuple}}(U_{i})=(m_{i},d_{i},H_{i},M_{i},S_{i})$ with the month, day, hour, minute, and second. From this tuple, we then compute a new time representation composed of the normalized cyclic month-hour pair $D_{i}^{\prime}=(\theta_{i},\phi_{i})$ using
@@ 3.3 Temporal Metric Learning
Since $\mathbf{p}_{i}$ and $\mathbf{q}_{i}$ are normalized, they have the same characteristics as probability mass functions. Thus, we define the image-time contrastive loss by computing the cross-entropy ($\mathrm{CE}$) between $\mathbf{p}_{i}$ and $\mathbf{q}_{i}$:
@@ 4 Experiments
Table 2: Geo-localization accuracy on Im2GPS3k & GWS15k datasets, reported on the ratio of samples that are correctly predicted under a distance threshold of 1 km radius.
@@ 4.2 Ablation Studies
In this section, we present additional experiments to evaluate the effectiveness of our proposed Temporal Metric Loss. Furthermore, in the supplementary material (Section 13), we explore alternative image backbones (i.e., DINOv2 [21], OpenCLIP [7]), different time encoders (i.e., Time2Vec [10], Circular Decomposition [17]), and various temporal resolutions to provide additional insights and justification for our design choices.
@@ 4.4 Compositional Image Retrieval
In this section, we evaluate GT-Loc’s capability for compositional image retrieval tasks, leveraging its unified multimodal embedding space. Specifically, we explore retrieving images based on joint queries consisting of both location and time information. Given query location $L^{Q}$ and time $T^{Q}$, we generate a multimodal representation by averaging their embeddings, inspired by multimodal retrieval methods proposed in prior works [31, 32].
@@ 13.2 Time representation
Motivated by the cyclical nature of time, Mac Aodha et al. 17 used circular decomposition to wrap the temporal input to their geographical prior encoder, resulting in similar embeddings for dates that are close to the start and end of the year, such as December $31^{st}$ and January $1^{st}$. To achieve this, for each dimension $l$ of the temporal input $\vec{x}$, they perform the mapping $[\sin(\pi x^{l}),\cos(\pi x^{l})]$, resulting in two numbers for each dimension.
@@ 14 Compositional image retrieval details
[…]
@@ 21 Limitations and reproducibility challenges of existing time prediction methods
[…]